OpenStudy (anonymous):

if the domain is {-3> or equal to x < or equal to}, state the range for f(x)^2 + 4 and g(x) = |x+1|

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

"{-3> or equal to x < or equal to}" seems all messed up

jimthompson5910 (jim_thompson5910):

also the same with "f(x)^2 + 4"

OpenStudy (anonymous):

\[-3\le x \le 3\]

OpenStudy (anonymous):

\[f(x)= x^{2} +4\]

jimthompson5910 (jim_thompson5910):

what you can do is go to this graphing calculator here https://www.desmos.com/calculator and type in each function. You would then place the domain at the end in curly braces example: x^2+4{-3 <= x <= 3} would plot f(x) based on that restriction

jimthompson5910 (jim_thompson5910):

then use the graph to read off the range

OpenStudy (anonymous):

we're not allowed to use graphing calculators

jimthompson5910 (jim_thompson5910):

ok plug in the x values of the domain into f(x) to get values that lie in the range of f(x) the range will span from the smallest output to the largest output

OpenStudy (anonymous):

oh so (-3)^2 + 4 and then (3) ^2 + 4

OpenStudy (anonymous):

isnt it just 13?

jimthompson5910 (jim_thompson5910):

that is the largest value in the range of f(x), what is the smallest?

OpenStudy (anonymous):

isnt that 13 too the - becomes positive 9 +4 so theyre both 13

jimthompson5910 (jim_thompson5910):

you have to plug in all the x values between -3 and 3

OpenStudy (anonymous):

ohh

jimthompson5910 (jim_thompson5910):

and see which y value is the smallest and which is the largest

OpenStudy (anonymous):

oh so 4

OpenStudy (anonymous):

\[4\le x \le 13\]

jimthompson5910 (jim_thompson5910):

yes so the range is \[\Large 4 \le y \le 13\]

jimthompson5910 (jim_thompson5910):

use y instead of x for the range

OpenStudy (anonymous):

oh right y yeah thanks!

jimthompson5910 (jim_thompson5910):

you'll do the same for g(x)

OpenStudy (anonymous):

so i got \[1\le y \le4\]

jimthompson5910 (jim_thompson5910):

there's a value smaller than 1

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

the lowest is 0+1

jimthompson5910 (jim_thompson5910):

did you try any negative x values to plug in?

OpenStudy (anonymous):

yeah but its absolute value so everythings positive

jimthompson5910 (jim_thompson5910):

what x values did you try so far

OpenStudy (anonymous):

oh i see what i did so the lowest is 0 right?

jimthompson5910 (jim_thompson5910):

yes so the range for g(x) is \[\Large 0 \le y \le 4\]

OpenStudy (anonymous):

yeah got it! thanks so much!

jimthompson5910 (jim_thompson5910):

yw